# Thermoelectric coefficients of two-flavor quark matter from the Kubo formalism

### v1

@@ Abstract
The hot quark matter created in heavy-ion collision experiments can exhibit strong temperature and chemical-potential gradients, which in turn can generate electric fields through thermoelectric effects. In this work, we investigate two relevant thermoelectric coefficients—the thermopower (Seebeck coefficient) and the Thomson coefficient—of two-flavor quark matter using the Kubo formalism and the Nambu–Jona-Lasinio model as an effective description of dense, finite-temperature QCD. The required two-point equilibrium correlation functions are evaluated using the Matsubara formalism of thermal field theory, applying a $1/N_{c}$ expansion to the relevant multi-loop Feynman diagrams. We employ previously derived quark spectral functions obtained from one–meson-exchange diagrams above the Mott transition temperature. Our numerical results show that both thermoelectric coefficients increase approximately linearly with temperature and decrease with increasing chemical potential. We also estimate the magnitude of the electric fields that can be generated in heavy-ion collisions by thermal gradients via the Seebeck effect.
@@ I Introduction
Thermoelectric phenomena in hot and dense QCD matter—most notably the Seebeck, Nernst, and related thermomagnetic transport coefficients—have attracted increasing interest as probes of non-equilibrium dynamics in systems with strong temperature gradients and magnetic fields; for recent reviews see Hattori:2022hyo ; Shovkovy:2025yvn . These coefficients offer sensitive diagnostics of quark–gluon plasma (QGP) properties, the evolution of electromagnetic fields, and the interplay between microscopic scattering processes, momentum-space anisotropies, chiral structure, and macroscopic hydrodynamic behavior. Their relevance extends from relativistic heavy-ion collisions, where intense short-lived electromagnetic fields and steep spatial gradients can induce substantial charge and heat transport, to astrophysical environments such as neutron stars, core-collapse supernovae, and binary-merger remnants, where dense quark matter, strong fields, and large thermodynamic gradients naturally occur. Over the past decade, a broad range of theoretical frameworks—including kinetic theory, effective field theory, and holography—have been employed to characterize these transport properties across both hadronic and partonic phases of strongly interacting matter.
@@
Strong-coupling studies based on holographic models provided some of the earliest insights into thermoelectric conductivities, with analyses such as Kim2015 offering benchmarks for Seebeck and Nernst coefficients at finite magnetic field. Complementary kinetic-theory calculations in hadronic matter followed, including studies of the Seebeck effect in hadron resonance gas models Bhatt2019 ; Das2020 . Thermoelectric response in QCD matter has since been explored in a variety of magnetic-field regimes. Works such as Dey2020 ; Dey2021 analyzed Seebeck and related coefficients in quantizing and weak magnetic fields, while the impact of collisional dynamics was studied in Kurian2021 using relaxation-time and BGK-type approaches. The role of plasma anisotropy was emphasized in Zhang2021 ; Rath2025 , which demonstrated that both Seebeck and Nernst coefficients are highly sensitive to momentum-space deformations in the QGP. More recent developments updated perturbative QCD treatments of the Seebeck effect Shaikh2025 , and analyses incorporating medium-dependent quark masses in anisotropic plasmas Khan2024 . Time-dependent magnetic fields — an essential feature of heavy-ion collisions — have been addressed in studies of both hadronic and partonic matter Kumar2025 ; Singh2024 ; Singh2025 , highlighting the dynamical interplay between evolving fields and induced thermoelectric currents.
@@
Effective model approaches have provided additional perspectives on transport in quark matter. For example, NJL-based studies such as Abhishek2020 examined thermoelectric coefficients using relaxation times derived from meson-exchange scattering, yielding a complementary quasiparticle description within strongly interacting matter. Two principal theoretical approaches are available for obtaining transport coefficients from an underlying microscopic description. For weakly interacting systems, kinetic theory based on the Boltzmann equation provides a natural framework for computing transport in terms of quasiparticle scattering processes. In contrast, strongly interacting quantum systems require a fully quantum-statistical treatment, where transport coefficients are related to appropriate equilibrium correlation functions via the Kubo formulas. The latter approach, grounded in the Liouville equation for the non-equilibrium density matrix, is applicable in principle to systems with strong interactions and collective dynamics.
@@
In this work, we employ the Kubo formalism to compute the thermal conductivity and thermoelectric power of hot quark matter. Specifically, thermoelectric power plays a crucial role when temperature gradients and electric fields coexist, as in the early stages of heavy-ion collisions. We adopt the Nambu–Jona-Lasinio (NJL) model, an effective chiral theory where quark–quark interactions are encoded through four-fermion contact terms. The non-equilibrium density matrix is constructed by perturbing the grand canonical ensemble and solving the Liouville equation to linear order in the applied electric field and temperature gradient. The relevant two-point correlation functions of electric and heat currents are evaluated in the Matsubara (imaginary-time) formalism, which, at leading order, are given by single-loop diagrams with fully dressed quark propagators, while vertex corrections are suppressed in $1/N_{c}$ counting scheme.
@@ II Derivation of the Kubo formulas
The Kubo formalism relates the transport coefficients of a statistical ensemble to certain equilibrium correlation functions, which in turn can be computed using equilibrium many-body techniques 1957JPSJ…12..570K ; 1957JPSJ…12.1203K ; Mahan . The thermodynamic state of a macroscopic quantum system can be described by means of the time-dependent density matrix $\rho(t)$. The density matrix (grand canonical distribution)
@@
applies to the equilibrium system with unperturbed Hamiltonian $\mathcal{H}_{0}$ and conserved particle number $\mathcal{N}$, where $\beta$ is the inverse temperature, $\mu$ is the chemical potential and $\Omega_{0}=-\ln\left({\rm Tr}e^{-\beta K_{0}}\right)$ is the grand thermodynamic potential. We will assume that the system is described by the equilibrium density matrix at the initial point of time, which we take at $t=$-$\infty$.
The derivation of the Kubo formulas for thermoelectric transport coefficients proceeds as follows. In a non-equilibrium state, the density matrix becomes time-dependent and obeys the Liouville equation, which in the Schrödinger representation reads
@@
where $\mathcal{H}$ is the Hamiltonian of the system including the external perturbation. In this representation, the operators acting on the quantum states of the system are time-independent; thus, the explicit time dependence of $\rho(t)$ arises solely from the time dependence of external fields and thermodynamic parameters. In what follows, we use the subscript $H$ to denote operators in the Heisenberg representation, while operators without a subscript are understood to be in the Schrödinger representation.
@@
One should keep in mind, however, that the local equilibrium density matrix (5) does not satisfy the Liouville equation (2), and therefore cannot describe irreversible processes. However, it can serve as a starting point (i.e., the zeroth-order approximation) to construct an exact density matrix at the required order in thermodynamic perturbations, such as external fields and thermal gradients. In this work, we are interested only in linear response of the system to thermoelectric perturbations. The derivation of the non-equilibrium density matrix up to the second order in perturbations in full relativistic framework was carried out recently in Refs. Harutyunyan2022 ; Harutyunyan2025 .
@@
Note that the local equilibrium density matrix $\rho_{l}$ does not commute with the Hamiltonian, because of the non-uniformity of temperature and chemical potential, as well as the presence of external electric field (in equilibrium we would have ${\cal H}_{0},\rho_{0}]=0$). Thus, it can be anticipated that the commutator ${\cal H},\rho_{l}]$ should be proportional to the gradients of these quantities and the electric field, as shown below. The detailed procedure of solving Eq. (8) is given in Appendix A. The result reads
@@
Next, we substitute the expressions for $V(t)$ and $W(t)$ given by Eqs. (10) and (11), assuming homogeneous electric and thermal fields with monochromatic time-dependence. Furthermore, we will assume that the perturbations are switched on adiabatically from the moment $t\to$-$\infty$, thus taking the perturbations in the form $E_{k}(\bm{r},t)=E_{k}e^{(s-i\omega)t}$, $\partial_{k}\beta(\bm{r},t)=\partial_{k}\beta e^{(s-i\omega)t}$ with $s\to+0$, where $E_{k},\partial_{k}\beta={\rm const}$. Then we obtain
@@
Now we express $\bm{h}$ in terms of $\bm{\nabla}T$ and $\bm{j}$ and define the thermoelectric power (thermopower) $Q=\alpha/\sigma=\gamma/T\sigma$, and thermal conductivity $\kappa=\tilde{\kappa}-TQ^{2}\sigma$ (the relation $\gamma=\alpha T$ follows from the symmetry property of the Green’s function 1957JPSJ…12..570K ; Harutyunyan2022 ; Harutyunyan2025 ). Finally, in the simultaneous presence of electric field $\bm{E}$ and thermal gradient $\bm{\nabla}T$, the statistical averages of the electrical and thermal currents in linear response theory can be written in the form
@@
Performing the temporal integrations in the expressions (19)–(22), the Kubo formulas for three thermoelectric coefficients can be derived, see App. B for details
@@
Eqs. (26)–(31) are known as Kubo formulas 1957JPSJ…12..570K ; 1957JPSJ…12.1203K ; Mahan . These equations apply to an arbitrary quantum statistical ensemble without restrictions on the strength of the couplings of the underlying theory. In the following, we will derive more specific expressions suitable for the NJL model with contact scalar and pseudoscalar couplings among quarks by applying the $1/N_{c}$ expansion to select the dominant diagrams contributing to the correlation functions.
The relations (25) lead to several thermoelectric effects, among which we will discuss the Seebeck effect and the Thomson effect (see, e.g., Refs Ziman1979 ; Landau8 ). The Seebeck effect refers to the generation of an electric field in a conducting material due to a temperature gradient in the absence of an electric current: $\bm{E}=Q\bm{\nabla}T$. The Seebeck coefficient, $Q$, is defined as the proportionality between $\bm{\nabla}T$ and $\bm{E}$ and is identical to the thermopower.
@@
${\cal\bm{Q}}=-{\rm div}\bm{h}+\bm{j}\cdot\bm{E}={\rm div}(\kappa\bm{\nabla}T)+\frac{\bm{j}^{2}}{\sigma}-T\bm{j}\bm{\nabla}Q,$ (32)
@@
Since our focus here is hot quark matter created in heavy-ion collisions, it makes sense to consider thermoelectric effects at constant baryon density, which is one-third of the quark number density $n$. Then the change in the thermopower $Q$ is related only to the temperature gradient, therefore the heat released due to the simultaneous presence of electric and heat fluxes (the Thomson effect) will be $\rho\bm{j}\bm{\nabla}T$, where
@@
is called the Thomson coefficient. Note that the Thomson heat changes sign when the direction of the current is reversed, so the Thomson coefficient can be either positive or negative. For $\rho>0$, heat is released when the current flows toward higher temperatures and absorbed when it flows in the opposite direction. For $\rho<0$, the situation is reversed.
@@ III Correlation functions in the NJL model
We consider two-flavor quark matter described by the NJL Lagrangian 1961PhRv..124..246N ; 1961PhRv..122..345N ; 1991PrPNP..27..195V ; 1992RvMP…64..649K ; 2005PhR…407..205B
@@
where $\psi=(u,d)^{T}$ is the iso-doublet quark field, $m_{0}=5.5$ MeV is the current quark mass, $G=10.1$ GeV-2 is the effective four-fermion coupling constant, and $\bm{\tau}$ is the vector of Pauli isospin matrices. This Lagrangian describes four-fermion scalar-isoscalar and pseudoscalar-isovector interactions between quarks with the corresponding bare vertices $\Gamma^{0}_{s}=1$ and $\Gamma^{0}_{ps}=i\bm{\tau}\gamma_{5}$. Four-fermion contact interactions make the model non-renormalizable; therefore, a 3-momentum cutoff ($p\leq\Lambda$) should be applied to regularize the ultraviolet divergences. The cutoff parameter $\Lambda$, the current quark masses $m_{0}$, and the four-fermion coupling $G$ constitute the free parameters of the model. Their values are chosen to reproduce experimental values of meson masses and the pion decay constant. Below, we will adopt the value $\Lambda=0.65$ GeV.
@@
and the quark-number and charge currents are defined as
@@
Next, we evaluate the correlation functions given by Eqs. (29)–(31) in thermal equilibrium using the imaginary-time Matsubara technique (see, e.g., Ref. Mahan for details). We mainly follow the derivations of Ref. Harutyunyan2017 , but consider also the thermopower and thermoelectric correction to the thermal conductivity, which were omitted there.
@@
where $T_{10}(\bm{r},\tau)$, $N_{1}(\bm{r},\tau)$, and $J_{1}(\bm{r},\tau)$ are obtained from $T_{10}(\bm{r},t)$, $N_{1}(\bm{r},t)$, and $J_{1}(\bm{r},t)$ via the Wick rotation $t\to-i\tau$, and ${\cal T}_{\tau}$ is the time-ordering operator for imaginary time $\tau$. In Eqs. (III) and (39) we substituted the operator of heat flux (112) and used the symmetry of the correlation function with respect to its arguments 1957JPSJ…12..570K ; Harutyunyan2017 . The required retarded correlation functions (29)–(31) can be obtained from their Matsubara counterparts (III)–(39) by an analytic continuation $i\omega_{n}\to\omega+i\delta$. Note that the transformation to imaginary time implies a change of the derivative $\partial_{0}\to i\partial_{\tau}$. Because $T_{\mu\nu}$, $N_{\mu}$ and $J_{\mu}$ are bosonic operators, the Matsubara frequencies assume even integer values $\omega_{n}=2\pi nT$, $n=0,\pm 1,\ldots$. The $T_{10}$ component of the energy-momentum tensor (35) in the Matsubara frequency space is given by
@@
Each of the correlation functions above can be represented as an infinite sum of diagrams shown schematically in Fig. 1 [here the operators $\hat{a}$ and $\hat{b}$ contain momentum and $\gamma$-matrices; their form for each correlation function should be specified be means of Eqs. (36) and (46)]. The propagators here are dressed, i.e., the dispersive effects are taken into account in their spectral width. In the $1/N_{c}$ expansion scheme, where $N_{c}$ is the number of colors, the diagrams in the first and second lines are of order ${\cal O}(N_{c}^{1})$, whereas the diagram in the third line is of order ${\cal O}(N_{c}^{0})$. Following the argument of Ref. Harutyunyan2017 (see also Refs. 1994PhRvC..49.3283Q ; 2008JPhG…35c5003I ; 2008EPJA…38…97A ; LW14 ; LKW15 ; 2016PhRvC..93d5205G ; LangDiss , we retain only the first-line diagram, as the others shown, along with higher-order diagrams in $G$, are subdominant.
@@
Substituting Eqs. (61)–(III) into Eqs. (III)–(39) and then into (26)–(28), we obtain the following expressions for thermoelectric transport coefficients
@@
where the coefficients $A_{s}=A_{s}(p,\varepsilon)$, $A_{0}=A_{0}(p,\varepsilon)$, $A_{v}=A_{v}(p,\varepsilon)$ depend only on $\bm{p}^{2}$ and $\varepsilon$ in isotropic medium. Note that the mass of the dressed quark $m$ is different from the current mass $m_{0}$ in the Lagrangian (34) because of the strong interactions 2005PhR…407..205B . Then we obtain for Eq. (57)

### v2

@@ Abstract
The hot quark matter created in heavy-ion collision experiments can exhibit strong temperature and chemical-potential gradients, which in turn can generate electric fields through thermoelectric effects. In this work, we investigate two relevant thermoelectric coefficients—the thermopower (Seebeck coefficient) and the Thomson coefficient—of two-flavor quark matter using the Kubo formalism and the Nambu–Jona-Lasinio model as an effective description of dense, finite-temperature QCD. The required two-point equilibrium correlation functions are evaluated using the Matsubara formalism of thermal field theory, applying a $1/N_{c}$ expansion to the relevant multiloop Feynman diagrams. We employ previously derived quark spectral functions obtained from one–meson-exchange diagrams above the Mott transition temperature. Our numerical results show that both thermoelectric coefficients increase approximately linearly with temperature and decrease with increasing chemical potential. We also estimate the magnitude of the electric fields that can be generated in heavy-ion collisions by thermal gradients via the Seebeck effect.
@@ I Introduction
Thermoelectric phenomena in hot and dense QCD matter—most notably the Seebeck, Nernst, and related thermomagnetic transport coefficients—have attracted increasing interest as probes of nonequilibrium dynamics in systems with strong temperature gradients and magnetic fields; for recent reviews see Hattori et al. (2022); Shovkovy and Ghosh (2025). These coefficients offer sensitive diagnostics of quark–gluon plasma (QGP) properties, the evolution of electromagnetic fields, and the interplay between microscopic scattering processes, momentum-space anisotropies, chiral structure, and macroscopic hydrodynamic behavior. Their relevance extends from relativistic heavy-ion collisions, where intense short-lived electromagnetic fields and steep spatial gradients can induce substantial charge and heat transport, to astrophysical environments such as neutron stars, core-collapse supernovae, and binary-merger remnants, where dense quark matter, strong fields, and large thermodynamic gradients naturally occur. Over the past decade, a broad range of theoretical frameworks—including kinetic theory, effective field theory, and holography—have been employed to characterize these transport properties across both hadronic and partonic phases of strongly interacting matter.
@@
Strong-coupling studies based on holographic models provided some of the earliest insights into thermoelectric conductivities, with analyses such as Kim et al. (2015) offering benchmarks for Seebeck and Nernst coefficients at finite magnetic field. Complementary kinetic-theory calculations in hadronic matter followed, including studies of the Seebeck effect in hadron resonance gas models Bhatt et al. (2019); Das et al. (2020). Thermoelectric response in QCD matter has since been explored in a variety of magnetic-field regimes. Works such as Dey and Patra (2020, 2021) analyzed Seebeck and related coefficients in quantizing and weak magnetic fields, while the impact of collisional dynamics was studied in Kurian (2021) using relaxation-time and BGK-type approaches. The role of plasma anisotropy was emphasized in Zhang et al. (2021); Rath and Neill (2025), which demonstrated that both Seebeck and Nernst coefficients are highly sensitive to momentum-space deformations in the QGP. More recent developments updated perturbative QCD treatments of the Seebeck effect Shaikh et al. (2025), and analyses incorporating medium-dependent quark masses in anisotropic plasmas Khan et al. (2024). Time-dependent magnetic fields–an essential feature of heavy-ion collisions–have been addressed in studies of both hadronic and partonic matter Kumar et al. (2025); Singh et al. (2024); Singh and Sahoo (2025), highlighting the dynamical interplay between evolving fields and induced thermoelectric currents.
@@
Effective model approaches have provided additional perspectives on transport in quark matter. For example, Nambu–Jona-Lasinio– (NJL) based studies such as Abhishek et al. (2022); Zhang et al. (2022) examined thermoelectric coefficients using relaxation times derived from meson-exchange scattering, yielding a complementary quasiparticle description within strongly interacting matter. Two principal theoretical approaches are available for obtaining transport coefficients from an underlying microscopic description. For weakly interacting systems, kinetic theory based on the Boltzmann equation provides a natural framework for computing transport in terms of quasiparticle scattering processes. In contrast, strongly interacting quantum systems require a fully quantum-statistical treatment, where transport coefficients are related to appropriate equilibrium correlation functions via the Kubo formulas. The latter approach, grounded in the Liouville equation for the nonequilibrium density matrix, is applicable in principle to systems with strong interactions and collective dynamics. These two approaches are complementary under certain conditions, as for example, in a weakly coupled relativistic scalar field theory, where sums of an infinite class of diagrams lead to linear integral equation, which are equivalent to a linearized Boltzmann equation Jeon (1995).
@@
In this work, we employ the Kubo formalism to compute the thermal conductivity and thermoelectric power of hot quark matter. Specifically, thermoelectric power plays a crucial role when temperature gradients and electric fields coexist, as in the early stages of heavy-ion collisions. We adopt the NJL model, an effective chiral theory where quark–quark interactions are encoded through four-fermion contact terms. The nonequilibrium density matrix is constructed by perturbing the grand canonical ensemble and solving the Liouville equation to linear order in the applied electric field and temperature gradient. The relevant two-point correlation functions of electric and heat currents are evaluated in the Matsubara (imaginary-time) formalism, which, at leading order, are given by single-loop diagrams with fully dressed quark propagators, while vertex corrections are suppressed in $1/N_{c}$ counting scheme. This type of approach has been applied in a broader context beyond the NJL model. For example, shear viscosity in field theories is dominated by one-loop diagrams with finite thermal widths, while higher-order loops are negligible Jeon (1993); Ghosh (2014). Similarly, electrical conductivities of baryon–meson systems have been computed using current–current correlators with thermal widths determined from in-medium scattering Ghosh (2017). We emphasize that, in the limit of vanishing widths, the correlators diverge and the transport coefficients vanish.
@@ II Derivation of the Kubo formulas
The Kubo formalism relates the transport coefficients of a statistical ensemble to certain equilibrium correlation functions, which in turn can be computed using equilibrium many-body techniques Kubo (1957); Kubo et al. (1957); Mahan (1990). The thermodynamic state of a macroscopic quantum system can be described by means of the time-dependent density matrix $\rho(t)$. The density matrix (grand canonical distribution)
@@
applies to the equilibrium system with unperturbed Hamiltonian $\mathcal{H}_{0}$ and conserved particle number $\mathcal{N}$, where $\beta$ is the inverse temperature, $\mu$ is the chemical potential, and $\Omega_{0}=-\ln\left({\rm Tr}e^{-\beta K_{0}}\right)$ is the grand thermodynamic potential. We will assume that the system is described by the equilibrium density matrix at the initial point of time, which we take at $t=$-$\infty$.
The derivation of the Kubo formulas for thermoelectric transport coefficients proceeds as follows. In a nonequilibrium state, the density matrix becomes time dependent and obeys the Liouville equation, which in the Schrödinger representation reads
@@
where $\mathcal{H}$ is the Hamiltonian of the system including the external perturbation. In this representation, the operators acting on the quantum states of the system are time independent; thus, the explicit time dependence of $\rho(t)$ arises solely from the time dependence of external fields and thermodynamic parameters. In what follows, we use the subscript $H$ to denote operators in the Heisenberg representation, while operators without a subscript are understood to be in the Schrödinger representation.
@@
One should keep in mind, however, that the local equilibrium density matrix (5) does not satisfy the Liouville equation (2) and therefore cannot describe irreversible processes. However, it can serve as a starting point (i.e., the zeroth-order approximation) to construct an exact density matrix at the required order in thermodynamic perturbations, such as external fields and thermal gradients. In this work, we are interested only in linear response of the system to thermoelectric perturbations. The derivation of the nonequilibrium density matrix up to the second order in perturbations in a full relativistic framework was carried out recently in Refs. Harutyunyan et al. (2022); Harutyunyan and Sedrakian (2025).
@@
Note that the local equilibrium density matrix $\rho_{l}$ does not commute with the Hamiltonian, because of the nonuniformity of temperature and chemical potential, as well as the presence of external electric field (in equilibrium we would have ${\cal H}_{0},\rho_{0}]=0$). Thus, it can be anticipated that the commutator ${\cal H},\rho_{l}]$ should be proportional to the gradients of these quantities and the electric field, as shown below. The detailed procedure of solving Eq. (8) is given in Appendix A. The result reads
@@
Next, we substitute the expressions for $V(t)$ and $W(t)$ given by Eqs. (10) and (11), assuming homogeneous electric and thermal fields with monochromatic time dependence. Furthermore, we will assume that the perturbations are switched on adiabatically from the moment $t\to$-$\infty$, thus taking the perturbations in the form $E_{k}(\bm{r},t)=E_{k}e^{(s-i\omega)t}$, $\partial_{k}\beta(\bm{r},t)=\partial_{k}\beta e^{(s-i\omega)t}$ with $s\to+0$, where $E_{k},\partial_{k}\beta={\rm const}$. Then we obtain
@@
Now we express $\bm{h}$ in terms of $\bm{\nabla}T$ and $\bm{j}$ and define the thermoelectric power (thermopower) $Q=\alpha/\sigma=\gamma/T\sigma$, and thermal conductivity $\kappa=\tilde{\kappa}-TQ^{2}\sigma$ (the relation $\gamma=\alpha T$ follows from the symmetry property of the Green’s function Kubo (1957); Harutyunyan et al. (2022); Harutyunyan and Sedrakian (2025)). Finally, in the simultaneous presence of electric field $\bm{E}$ and thermal gradient $\bm{\nabla}T$, the statistical averages of the electrical and thermal currents in linear response theory can be written in the form
@@
Performing the temporal integrations in the expressions (19)–(22), the Kubo formulas for three thermoelectric coefficients can be derived, see Appendix B for details,
@@
Equations (26)–(31) are known as Kubo formulas Kubo (1957); Kubo et al. (1957); Mahan (1990). These equations apply to an arbitrary quantum-statistical ensemble without restrictions on the strength of the couplings of the underlying theory. In the following, we will derive more specific expressions suitable for the NJL model with contact scalar and pseudoscalar couplings among quarks by applying the $1/N_{c}$ expansion to select the dominant diagrams contributing to the correlation functions.
The relations (25) lead to several thermoelectric effects, among which we will discuss the Seebeck effect and the Thomson effect (see, e.g., Refs Ziman (1979); Landau and Lifshitz (1960)). The Seebeck effect refers to the generation of an electric field in a conducting material due to a temperature gradient in the absence of an electric current: $\bm{E}=Q\bm{\nabla}T$. The “Seebeck coefficient” $Q$ is defined as the proportionality between $\bm{\nabla}T$ and $\bm{E}$ and is identical to the thermopower.
@@
${\cal\bm{Q}}=-{\rm div}\bm{h}+\bm{j}\cdot\bm{E}={\rm div}(\kappa\bm{\nabla}T)+\frac{\bm{j}^{2}}{\sigma}-T\bm{j}\cdot\bm{\nabla}Q,$ (32)
@@
Since our focus here is hot quark matter created in heavy-ion collisions, it makes sense to consider thermoelectric effects at constant baryon density, which is one-third of the quark number density $n$. Then the change in the thermopower $Q$ is related only to the temperature gradient, therefore the heat released due to the simultaneous presence of electric and heat fluxes (the Thomson effect) will be $\rho\bm{j}\cdot\bm{\nabla}T$, where
@@
is called the “Thomson coefficient”. Note that the Thomson heat changes sign when the direction of the current is reversed, so the Thomson coefficient can be either positive or negative. For $\rho>0$, heat is released when the current flows toward higher temperatures and absorbed when it flows in the opposite direction. For $\rho<0$, the situation is reversed.
@@ III Correlation functions in the NJL model
We consider two-flavor quark matter described by the NJL Lagrangian Nambu and Jona-Lasinio (1961b, a); Vogl and Weise (1991); Klevansky (1992); Buballa (2005)
@@
where $\psi=(u,d)^{T}$ is the isodoublet quark field, $m_{0}=5.5$ MeV is the current quark mass, $G=10.1$ GeV-2 is the effective four-fermion coupling constant, and $\bm{\tau}$ is the vector of Pauli isospin matrices. This Lagrangian describes four-fermion scalar-isoscalar and pseudoscalar-isovector interactions between quarks with the corresponding bare vertices $\Gamma^{0}_{s}=1$ and $\Gamma^{0}_{ps}=i\bm{\tau}\gamma_{5}$. Four-fermion contact interactions make the model nonrenormalizable; therefore, a 3-momentum cutoff ($p\leq\Lambda$) should be applied to regularize the ultraviolet divergences. The cutoff parameter $\Lambda$, the current quark masses $m_{0}$, and the four-fermion coupling $G$ constitute the free parameters of the model. Their values are chosen to reproduce experimental values of meson masses and the pion decay constant. Below, we will adopt the value $\Lambda=0.65$ GeV.
@@
and the quark number and charge currents are defined as
@@
Next, we evaluate the correlation functions given by Eqs. (29)–(31) in thermal equilibrium using the imaginary-time Matsubara technique (see, e.g., Ref. Mahan (1990) for details). We mainly follow the derivations of Ref. Harutyunyan et al. (2017), but consider also the thermopower and thermoelectric correction to the thermal conductivity, which were omitted there.
@@
where $T_{10}(\bm{r},\tau)$, $N_{1}(\bm{r},\tau)$, and $J_{1}(\bm{r},\tau)$ are obtained from $T_{10}(\bm{r},t)$, $N_{1}(\bm{r},t)$, and $J_{1}(\bm{r},t)$ via the Wick rotation $t\to-i\tau$, and ${\cal T}_{\tau}$ is the time-ordering operator for imaginary time $\tau$. In Eqs. (III) and (39) we substituted the operator of heat flux (113) and used the symmetry of the correlation function with respect to its arguments Kubo (1957); Harutyunyan et al. (2017). The required retarded correlation functions (29)–(31) can be obtained from their Matsubara counterparts (III) and (39) by an analytic continuation $i\omega_{n}\to\omega+i\delta$. Note that the transformation to imaginary time implies a change of the derivative $\partial_{0}\to i\partial_{\tau}$. Because $T_{\mu\nu}$, $N_{\mu}$, and $J_{\mu}$ are bosonic operators, the Matsubara frequencies assume even integer values $\omega_{n}=2\pi nT$, $n=0,\pm 1,\ldots$. The $T_{10}$ component of the energy-momentum tensor (35) in the Matsubara frequency space is given by
@@
Each of the correlation functions above can be represented as an infinite sum of diagrams shown schematically in Fig. 1 [here the operators $\hat{a}$ and $\hat{b}$ contain momentum and $\gamma$ matrices; their form for each correlation function should be specified be means of Eqs. (36) and (46)]. The propagators here are dressed, i.e., the dispersive effects are taken into account in their spectral width. In the $1/N_{c}$ expansion scheme, where $N_{c}$ is the number of colors, the diagrams in the first and second lines are of order ${\cal O}(N_{c}^{1})$, whereas the diagram in the third line is of order ${\cal O}(N_{c}^{0})$. Following the argument of Ref. Harutyunyan et al. (2017) (see also Refs. Quack and Klevansky (1994); Iwasaki et al. (2008); Alberico et al. (2008); Lang and Weise (2014); Lang et al. (2015); Ghosh et al. (2016); Lang (2015), we retain only the first-line diagram, as the others shown, along with higher-order diagrams in $G$, are subdominant.
@@
Substituting Eqs. (61)–(III) into Eqs. (III) and (39) and then into (26)–(28), we obtain the following expressions for thermoelectric transport coefficients:
@@
where the coefficients $A_{s}=A_{s}(p,\varepsilon)$, $A_{0}=A_{0}(p,\varepsilon)$, $A_{v}=A_{v}(p,\varepsilon)$ depend only on $\bm{p}^{2}$ and $\varepsilon$ in isotropic medium. Note that the mass of the dressed quark $m$ is different from the current mass $m_{0}$ in the Lagrangian (34) because of the strong interactions Buballa (2005). Then we obtain for Eq. (57)